Provided that z(0) = 10,

Solve the differential equation of: dz/dx = -z

Secondly, solve the differential equation of: dz/dx = z^2

Now, state which of these is a linear differentiable equation?

State which has solutions for every x greater than or equal to zero and provide explanation.
Answer #1

Z (0)=10 dz = [da Inz = -utc (n(10) = -otc C=1 (n = anti (dz =de -1/2 = 4tc 2 = - the